# Strong-lensing cosmography using third-generation gravitational-wave detectors

### v1

@@ Abstract
We present a detailed exposition of a statistical method for estimating cosmological parameters from the observation of a large number of strongly lensed binary-black-hole (BBH) mergers observable by third-generation gravitational-wave (GW) detectors. This method, first presented in [1], compares the observed number of strongly lensed GW events and their time delay distribution (between lensed images) with observed events to infer cosmological parameters. We show that the precision of the estimation of the cosmological parameters does not have a strong dependance on the assumed BBH redshift distribution model. An incorrect model for the distribution of lens properties and redshifts can lead to a biased cosmological inference. However, Bayesian model selection can assist in selecting the right model from a set of available parametric models for the distribution of lens redshift and properties. We also present a way to incorporate the effect of contamination in the data due to the limited efficiency of lensing identification methods, so that it will not bias the cosmological inference.
  • May 28, 2024
@@ 1 Introduction
At high redshifts ($z\gtrsim 0.5$), it is difficult to observe EM counterparts of compact binary mergers. Additionally, BBH mergers, which can be seen out to very high redshifts are generally not expected to produce EM counterparts. However, even in the absence of EM counterparts, statistical correlation of the mergers with large scale structure will provide some way of inferring cosmological parameters [9, 10, 11, 12, 13]. Here the Hubble diagram will contain imprints of other cosmological parameters such as the dark matter energy density and the equation of state of the dark energy. Thus they will provide strong tests of cosmological models, such as the currently favoured $\Lambda$CDM model [14].
We are unable to measure cosmological redshifts from BBH observations because the masses are degenerate with redshift in the GW signal, thanks to the absence of a preferred mass scale in GR. However, non-gravity physics in compact objects and their progenitors could introduce particular mass scales in compact binaries. For example, neutron stars have a maximum mass that depends on the nuclear equation of state [15] and the BH mass distribution can have some features due to pair instability supernovae [16, 17, 18, 19]. If such features are well understood (see, e.g. [20, 21] for some caveats), they will enable redshifts measurements from compact binaries purely using GW observations [22, 23, 24, 25, 26, 27]. Such spectral sirens will also enable GW cosmography.
@@
In this paper, we discuss a novel new cosmological probe that GW observations will enable, making use of strongly lensed GWs. GR predicts that intervening massive objects such as galaxies and clusters between the GW source and the observer will deflect the GWs through the phenomenon of gravitational lensing. If the lensing objects are sufficiently massive and compact, and lie sufficiently close to the line of sight to the source, they can create multiple images of the source. This phenomenon, called strong gravitational lensing, is routinely observed in EM observations of galaxies, clusters, quasars, etc. The same objects should strongly lens GW signals as well. The precise fraction of strongly lensed GW sources will depend on the distribution of the GW sources, galaxies and clusters that act as lenses as well as the cosmological model and parameters. According to recent calculations, this fraction is $\sim 0.01-0.05\%$ for current generation detectors [32, 33] and $\sim 0.1-1\%$ for the XG detectors [1, 34, 35, 36, 37, 38, 39, 40, 41]. Since XG detectors are expected to detect millions of compact binary mergers during their operation, they will detect thousands or tens of thousands of strongly lensed GWs.
@@
The idea of the new method is that, the precise number of strongly lensed GW events and the distribution of their lensing time delay depends on the cosmological model and parameters, apart from the distribution of GW sources and lenses [1]. If the latter two are known, then this will provide a new means of probing the cosmological parameters. The distribution of GW sources can be accurately determined by the large number of un-lensed signals that will dominate the data (see, e.g., [42]). The knowledge of the distribution of large gravitational lenses should come from cosmological simulations and EM observations  [^1] See [43] for a complementary approach to constrain the distribution of the GW sources and lenses from the lensing rate and time delay distribution, assuming a cosmology.. However, we will also show that the cosmological parameters and the model of the distribution of the lenses have sufficiently different imprints on the number of strong lensed events and their time delays, so that it is possible to disentangle these effects to a very good extent.
@@
Our statistical method, based on the observation of a large number of strongly lensed BBH events, does not require the presence EM counterparts to the mergers. A complementary approach, presented in [44, 45, 46], makes use of a much smaller number of strongly mergers having an EM counterparts. The concurrent observation of lensed EM and GW signals will allow us to measure cosmological parameters from even a single event. However, due to the limited observing horizon of EM telescopes, such observations will be able to probe only low-redshift ($z\lesssim 0.5$) cosmology [^2] Combing GW lensing observations with lenses detected in EM surveys (e.g., Euclid, CSST and JWST) could probe cosmology at moderate redshifts ($z\lesssim 2$)  [41, 47]. However, the exact method for this approach is yet to be developed..
@@
This paper is organised as follows: In section 2 we briefly review the Bayesian method that we use to constrain the cosmological parameters from the observation of a population of strongly lensed GW signals. This method was first presented in [1]. In section 3, we forecast the expected constraints on cosmological parameters assuming various astrophysical models for the redshift distribution of BBH mergers. In the remaining sections we investigate the various sources of systematic errors in our analysis, and show that it is possible to bring them sufficiently small that interesting measurements are possible in future. In particular, in section 4, we investigate how our inaccurate understanding of the lens distributions will bias our cosmological inference. In section 5 we will explore how we can deal with contaminated data – that is, we develop a formalism deal with the presence of a small number of unlensed GW signals in our lensing data, that are misidentified as lensed events due to our limited ability to distinguish between lensed and unlensed GW signals.
@@ 2.1 Bayesian inference of cosmological parameters
where $p\left(\vec{\Omega}~{}|~{}\mathcal{M}\right)$ is the prior distribution of $\vec{\Omega}$ given some model $\mathcal{M}$ while $p\left(N,\{\Delta t_{i}\}~{}|~{}\vec{\Omega},T_{\mathrm{obs}},\mathcal{M}\ \right)$ is the likelihood of observing $N$ lensed events with time delays $\{\Delta t_{i}\}_{i=1}^{N}$ given the set of cosmological parameters $\vec{\Omega}$ and model $\mathcal{M}$. The normalization constant ${p\left(N,\{\Delta t_{i}\}~{}|~{}\mathcal{M}\right)}$ is the evidence of the assumed model $\mathcal{M}$.
@@
Above, $p(\Delta t_{i}~{}|~{}\vec{\Omega},T_{\mathrm{obs}})$, can be thought of as “model” time-delay distribution $p(\Delta t~{}|~{}\vec{\Omega},T_{\mathrm{obs}})$ evaluated at the measured $\Delta t_{i}$ of a lensed merger. The shape of the model distribution is governed by the cosmological parameters $\vec{\Omega}$. The model distribution $p(\Delta t~{}|~{}\vec{\Omega},T_{\mathrm{obs}})$ is obtained from the expected (intrinsic) time delay distribution $p(\Delta t~{}|~{}\vec{\Omega})$, after applying the condition that we can not observe the time delays which are greater than the observation time $T_{\mathrm{obs}}$:
$p\left(\Delta t~{}|~{}\vec{\Omega},T_{\mathrm{obs}}\right)\propto p\left(\Delta t~{}|~{}\vec{\Omega}\right){\left(T_{\mathrm{obs}}-\Delta t\right)}\,\Theta(T_{\mathrm{obs}}-\Delta t),\$ (6)
@@ 2.2.1 Expected number of lensed events:
Above, $R$ is the BBH detection rate, $p_{b}(z_{s}|\vec{\Omega})$ is the redshift distribution (probability density) of merging binaries and $P_{\ell}(z_{s}|\vec{\Omega})$ is the strong lensing probability for the source redshift $z_{s}$. Here we assume that the GW detectors are able to detect all the merging binaries out to $z_{\mathrm{max}}$. For XG detectors, this is a good assumption for the $z_{\mathrm{max}}$ values that we use  [^3] The $z_{\mathrm{max}}$ predicted by a source population model (e.g., [48]) assumes the standard cosmology $\vec{\Omega}_{\mathrm{true}}$. When we consider other values of $\vec{\Omega}$, we rescale $z_{\mathrm{max}}$ appropriately.. $\mathcal{S}(T_{\mathrm{obs}})$ denotes the selection effects due to the finite observing time
@@
We expect that rate $R$ of the BBH mergers and their redshift distribution $p_{b}(z_{s}|\vec{\Omega})$ will be accurately measured from the large ($\sim 10^{6}$) number of unlensed events that will dominate the data [^4] What we measure from GW observations is the distribution $p_{b}(d_{L})$ of luminosity distance of the sources, which can then be converted into a redshift distribution $p_{b}(z_{s}|\vec{\Omega})$ assuming a set of cosmological parameters $\vec{\Omega}$. For the forecast analysis presented in this paper, we create a luminosity distance distribution from a redshift distribution model, assuming standard cosmological parameters $\vec{\Omega}_{\mathrm{true}}$. By varying $\vec{\Omega}$, we can then obtain different redshift distributions which we use to model the time delay distribution for that specific value of $\vec{\Omega}$.. However, these quantities are currently poorly known. To forecast the expected precision in measuring the cosmological parameters, we take different theoretical models of $p_{b}(z_{s}|\vec{\Omega})$ assuming a BBH detection rate of $R=5\times 10^{5}$ per year. The corresponding forecasts are described in section 3.
To compute the expected number of lensed events using equation (7), we also need to know the probability $P_{\ell}(~{}z_{s}|\vec{\Omega})$ that a source redshift $z_{s}$ is strongly lensed. This will depend on the distribution of lenses as well as cosmological parameters. We assume that the lenses are modelled by the SIS model. Multiple images are produced when the projected location of the source in the lens plane is within the Einstein radius of the lens, given by
$r_{E}(\sigma,z_{\ell},z_{s},\vec{\Omega})=4\pi\left(\frac{\sigma}{c}\right)^{4}\frac{D_{\ell}(z_{\ell},\vec{\Omega})~{}D_{\ell s}(z_{\ell},z_{s},\vec{\Omega})}{D_{s}(z_{s},\vec{\Omega})},$ (9)
@@
where the first term denotes the comoving number density of dark matter halos with mass $M$ at redshift $z_{\ell}$ predicted by the cosmological model with parameters $\vec{\Omega}$. We consider several models of the halo mass function (HMF) calibrated to cosmological simulations. The second term is a Jacobian to convert the distribution of the halo mass to that of the dispersion velocity. To compute the Jacobian, we assume halos with spherical boundaries that contain an average density of $\rho$. Thus,
@@
We also need to use some minimum and maximum cutoff for $\sigma$ to compute the total optical depth defined in equation (10). The natural choices are $\sigma_{\mathrm{min}}=\sigma(M_{\mathrm{min}})$ and $\sigma_{\mathrm{max}}=\sigma(M_{\mathrm{max}})$. We assume $M_{\mathrm{min}}=10^{8}\;M_{\odot}$ and $M_{\mathrm{max}}=10^{14}\;M_{\odot}$, since this the mass range of validity of most of the HMF models that we use.
It is now easy to see from equation (12), why the cosmological parameters affect the lensing optical depth and hence the number of detected lensed events. The first term describes a purely geometrical effect of how the comoving volume at a given redshift varies with a change in cosmological parameters. The second term describes the change in the distribution of lenses due to changes in the structure formation. Third shows how the fractional area covered by lenses at a given redshift varies due to the geometric effect.
@@ 2.2.2 Expected distribution of lensing time delays:
  • The redshift distribution of GW sources: We expect that this can be measured with sufficient precision from the large number $\sim 10^{6}$ of unlensed events that will dominate the data (see, e.g., [42]). In section 3 we forecast the prospective constraints on cosmological parameters assuming various theoretical models of the source redshift distribution (figure 1).
@@ 3 Effect of source distribution models
Figure 1: Left: Different models for the distributions of redshifts for BBH mergers. These includes prediction from population synthesis studies by Dominik [48] and Belczynski [50, 51], as well as uniform in comoving volume. Other models are based on Madau-Dickinson star formation rate [52] and consider different distributions of the time delay between formation and merger. Middle: The distributions of time delay between two lensed images of a source for different models of redshift distribution of BBH mergers shown in the left panel. Dashed lines represent the actual time delay distributions $p(\Delta t~{}|~{}\vec{\Omega})$ while solid lines represent time delay distributions $p(\Delta t~{}|~{}\vec{\Omega},T_{\mathrm{obs}})$ that will be observable in $10\;\mathrm{yrs}$. Time delay distributions are calculated for $\vec{\Omega}=\{\Omega_{m}=0.316,\;H_{0}=67.3\}$. Right: Posteriors (95% credible region) on cosmological parameters $\left(\Omega_{m},H_{0}\right)$ for different models of redshift distributions. This is done considering a BBH merger rate of $5\times 10^{5}\;\mathrm{yr}^{-1}$ with observation time period of $10\;\mathrm{yrs}$. Dashed cross represents true cosmological parameters.
The left panel of figure 1 displays the redshift distributions of BBH mergers which we consider. We consider a “true” cosmology $\vec{\Omega}_{\mathrm{true}}=\{\Omega_{m}=0.316,\;H_{0}=67.3\}$. The true distribution of lens redshift and parameters are calculated using a HMF model as described by [53], implemented in HMFcalc package [54] as described in section 2.2. We assume a merger rate of $R=5\times 10^{5}\;\mathrm{yr}^{-1}$ for BBHs, with an observation time period $T_{\mathrm{obs}}=10\;\mathrm{yrs}$.
@@
Using the method outlined in section 2.1, we compute the posteriors on the cosmological parameters $\vec{\Omega}$ from the different observing scenarios corresponding to the different source redshift distributions. We also assume that the redshift distribution of the sources is measured with sufficient precision from the observation of unlensed events. Posteriors for for the cosmological parameters are shown in the right panel of figure 1. When the source populations extend to high redshift (e.g., Dominik, Uniform), the precision is better in comparison to populations where the merger redshift does not have support at high redshift (e.g., Belczynski, flatlog, etc.). This is because the lensing optical depth increases with redshift, and therefore, populations that extend to high redshift are expected to have a larger number of lensed events than the populations that don’t extend to high redshift. The expected 90% credible intervals in the posteriors of the cosmological parameters are summarised in table 1, for different source redshift distribution models.

### v2

@@ Abstract
We present a detailed exposition of a statistical method for estimating cosmological parameters from the observation of a large number of strongly lensed binary-black-hole (BBH) mergers observable by next (third) generation (XG) gravitational-wave (GW) detectors. This method, first presented in Jana (2023 Phys. Rev. Lett. 130 261401), compares the observed number of strongly lensed GW events and their time delay distribution (between lensed images) with observed events to infer cosmological parameters. We show that the precision of the estimation of the cosmological parameters does not have a strong dependance on the assumed BBH redshift distribution model. Using the large number of unlensed mergers, XG detectors are expected to measure the BBH redshift distribution with sufficient precision for the cosmological inference. However, a biased inference of the BBH redshift distribution will bias the estimation of cosmological parameters. An incorrect model for the distribution of lens properties can also lead to a biased cosmological inference. However, Bayesian model selection can assist in selecting the right model from a set of available parametric models for the lens distribution. We also present a way to incorporate the effect of contamination in the data due to the limited efficiency of lensing identification methods, so that it will not bias the cosmological inference.
  • November 19, 2024
@@ 1 Introduction
At high redshifts ($z\gtrsim 0.5$), it is difficult to observe EM counterparts of compact binary mergers. Additionally, BBH mergers, which can be seen out to very high redshifts, are generally not expected to produce EM counterparts. However, even in the absence of EM counterparts, statistical correlation of the mergers with large scale structure will provide some way of inferring cosmological parameters [8, 9, 10, 11, 12]. Here the Hubble diagram will contain imprints of other cosmological parameters such as the dark matter energy density and the equation of state of the dark energy. Thus they will provide strong tests of cosmological models, such as the currently favoured $\Lambda$CDM model [13].
We are unable to measure cosmological redshifts from BBH observations because the masses are degenerate with redshift in the GW signal, thanks to the absence of a preferred mass scale in GR. However, non-gravity physics in compact objects and their progenitors could introduce particular mass scales in compact binaries. For example, neutron stars have a maximum mass that depends on the nuclear equation of state [14] and the BH mass distribution can have some features due to pair instability supernovae [15, 16, 17, 18]. If such features are well understood (see e.g. [19, 20] for some caveats), they will enable redshifts measurements from compact binaries purely using GW observations [21, 22, 23, 24, 25, 26]. Such spectral sirens will also enable GW cosmography.
@@
In this paper, we discuss a new cosmological probe that GW observations will enable, making use of strongly lensed GWs. GR predicts that intervening massive objects such as galaxies and clusters between the GW source and the observer will deflect the GWs through the phenomenon of gravitational lensing. If the lensing objects are sufficiently massive and compact, and lie sufficiently close to the line of sight to the source, they can create multiple images of the source. This phenomenon, called strong gravitational lensing, is routinely observed in EM observations of galaxies, clusters, quasars, etc. The same objects should strongly lens GW signals as well. The precise fraction of strongly lensed GW sources will depend on the distribution of the GW sources, galaxies and clusters that act as lenses as well as the cosmological model and parameters. According to recent calculations, this fraction is $\sim 0.01-0.05\%$ for current generation detectors [31, 32] and $\sim 0.1-1\%$ for the XG detectors [33, 34, 35, 36, 37, 38, 39, 40, 41]. Since XG detectors are expected to detect millions of compact binary mergers during their operation, they will detect thousands or tens of thousands of strongly lensed GWs.
@@
The idea of the new method is that the precise number of strongly lensed GW events and the distribution of their lensing time delay depends on the cosmological model and parameters, apart from the distribution of GW sources and lenses [33]. If the latter two are known, then this will provide a new means of probing the cosmological parameters. The distribution of GW sources can be accurately determined by the large number of un-lensed signals that will dominate the data (see, e.g., [42]). The knowledge of the distribution of large gravitational lenses should come from cosmological simulations and EM observations  [^2] See [34] for a complementary approach to constrain the distribution of the GW sources and lenses from the lensing rate and time delay distribution, assuming a cosmology.. However, we will also show that the cosmological parameters and the model of the distribution of the lenses have sufficiently different imprints on the number of strong lensed events and their time delays, so that it is possible to disentangle these effects to a very good extent.
@@
Our statistical method, based on the observation of a large number of strongly lensed BBH events, does not require the presence of EM counterparts to the mergers. A complementary approach, presented in [43, 44, 45], makes use of a much smaller number of strongly lensed mergers having an EM counterpart. The concurrent observation of lensed EM and GW signals will allow us to measure cosmological parameters from even a single event. However, due to the limited observing horizon of EM telescopes, such observations will be able to probe only low-redshift ($z\lesssim 0.5$) cosmology [^3] Combing GW lensing observations with lenses detected in EM surveys (e.g., Euclid, CSST and JWST) could probe cosmology at moderate redshifts ($z\lesssim 2$)  [41, 46]. However, the exact method for this approach is yet to be developed..
@@
This paper is organised as follows: In section 2 we briefly review the Bayesian method that we use to constrain the cosmological parameters from the observation of a population of strongly lensed GW signals. This method was first presented in [33]. In section 3, we forecast the expected constraints on cosmological parameters assuming various astrophysical models for the redshift distribution of BBH mergers. In the remaining sections we investigate the various sources of systematic errors in our analysis, and show that it is possible to bring them sufficiently small so that interesting measurements are possible in the future. In particular, in section 4, we explore how the errors in the measurement of the luminosity distance of individual GW events will limit our ability to reconstruct the true redshift distribution of GW sources, hence biasing our cosmological inference. In section 5, we investigate how our inaccurate understanding of the lens distributions will bias our cosmological inference. In section 6 we will explore how we can deal with contaminated data – that is, we develop a formalism deal with the presence of a small number of unlensed GW signals in our lensing data, that are misidentified as lensed events due to our limited ability to distinguish between lensed and unlensed GW signals.
@@ 2.1 Bayesian inference of cosmological parameters
where $p\left(\vec{\Omega}~{}|~{}\mathcal{M}\right)$ is the prior distribution of $\vec{\Omega}$ given some model $\mathcal{M}$ while $p\left(N,\{\Delta t_{i}\}~{}|~{}\vec{\Omega},T_{\mathrm{obs}},\mathcal{M}\ \right)$ is the likelihood of observing $N$ lensed events with time delays $\{\Delta t_{i}\}_{i=1}^{N}$ given the set of cosmological parameters $\vec{\Omega}$ and model $\mathcal{M}$. The normalisation constant ${p\left(N,\{\Delta t_{i}\}~{}|~{}\mathcal{M}\right)}$ is the evidence of the assumed model $\mathcal{M}$.
@@
Above, $p(\Delta t_{i}~{}|~{}\vec{\Omega},T_{\mathrm{obs}})$, can be thought of as a “model” time-delay distribution $p(\Delta t~{}|~{}\vec{\Omega},T_{\mathrm{obs}})$ evaluated at the measured $\Delta t_{i}$ of a lensed merger. The shape of the model distribution is governed by the cosmological parameters $\vec{\Omega}$. The model distribution $p(\Delta t~{}|~{}\vec{\Omega},T_{\mathrm{obs}})$ is obtained from the expected (intrinsic) time delay distribution $p(\Delta t~{}|~{}\vec{\Omega})$, after applying the condition that we can not observe the time delays which are greater than the observation time $T_{\mathrm{obs}}$:
$p\left(\Delta t~{}|~{}\vec{\Omega},T_{\mathrm{obs}}\right)\propto p\left(\Delta t~{}|~{}\vec{\Omega}\right){\left(T_{\mathrm{obs}}-\Delta t\right)}\,\Theta(T_{\mathrm{obs}}-\Delta t),$ (6)
@@ 2.2.1 Expected number of lensed events:
Above, $R$ is the BBH detection rate, $p_{b}(z_{s}|\vec{\Omega})$ is the redshift distribution (probability density) of merging binaries and $P_{\ell}(z_{s}|\vec{\Omega})$ is the strong lensing probability for the source redshift $z_{s}$. Here we assume that the GW detectors are able to detect all the merging binaries out to $z_{\mathrm{max}}$. For XG detectors, this is a good assumption for the $z_{\mathrm{max}}$ values that we use  [^4] The $z_{\mathrm{max}}$ predicted by a source population model (e.g., [47]) assumes the standard cosmology $\vec{\Omega}_{\mathrm{true}}$. For the population models that we consider, $z_{\mathrm{max}}\simeq 20$. When we consider other values of $\vec{\Omega}$, we rescale $z_{\mathrm{max}}$ appropriately.. $\mathcal{S}(T_{\mathrm{obs}})$ denotes the selection effects due to the finite observing time
@@
We expect that the rate $R$ of the BBH mergers and their redshift distribution $p_{b}(z_{s}|\vec{\Omega})$ will be accurately measured from the large ($\sim 10^{6}$) number of unlensed events that will dominate the data [^5] What we measure from GW observations is the distribution $p_{b}(d_{L})$ of luminosity distance of the sources, which can then be converted into a redshift distribution $p_{b}(z_{s}|\vec{\Omega})$ assuming a set of cosmological parameters $\vec{\Omega}$. For the forecast analysis presented in this paper, we create a luminosity distance distribution from a redshift distribution model, assuming standard cosmological parameters $\vec{\Omega}_{\mathrm{true}}$. By varying $\vec{\Omega}$, we can then obtain different redshift distributions which we use to model the time delay distribution for that specific value of $\vec{\Omega}$.. However, these quantities are currently poorly known. To forecast the expected precision in measuring the cosmological parameters, we take different theoretical models of $p_{b}(z_{s}|\vec{\Omega})$ assuming a BBH detection rate of $R=5\times 10^{5}$ per year. The corresponding forecasts are described in section 3. In section 4 we explore the effect of the GW measurement errors in the reconstruction of $p_{b}(z_{s}|\vec{\Omega})$ and hence on the inference of cosmological parameters.
To compute the expected number of lensed events using equation (7), we also need to know the probability $P_{\ell}(~{}z_{s}|\vec{\Omega})$ that a source at redshift $z_{s}$ is strongly lensed. This will depend on the distribution of lenses as well as cosmological parameters. We assume that the lenses are modelled by the SIS model. Multiple images are produced when the projected location of the source in the lens plane is within the Einstein radius of the lens, given by
$r_{E}(\sigma,z_{\ell},z_{s},\vec{\Omega})=4\pi\left(\frac{\sigma}{c}\right)^{2}\frac{D_{\ell}(z_{\ell},\vec{\Omega})~{}D_{\ell s}(z_{\ell},z_{s},\vec{\Omega})}{D_{s}(z_{s},\vec{\Omega})},$ (9)
@@
where the first term denotes the comoving number density of dark matter halos with mass $M$ at redshift $z_{\ell}$ predicted by the cosmological model with parameters $\vec{\Omega}$. We consider several models of the halo mass function (HMF) calibrated to cosmological simulations. The second term is a Jacobian to convert the distribution of the halo mass to that of the dispersion velocity. To compute the Jacobian, we assume that the halos are spherically symmetric and virialised, with uniform density $\rho$ and radius $R$. Thus,
@@
We also need to use some minimum and maximum cutoff for $\sigma$ to compute the total optical depth defined in equation (10). The natural choices are $\sigma_{\mathrm{min}}=\sigma(M_{\mathrm{min}})$ and $\sigma_{\mathrm{max}}=\sigma(M_{\mathrm{max}})$. We assume $M_{\mathrm{min}}=10^{8}\;M_{\odot}$ and $M_{\mathrm{max}}\simeq 10^{15}\;M_{\odot}$, since this the mass range of validity of most of the HMF models that we use.
It is now easy to see from equation (11) why the cosmological parameters affect the lensing optical depth and hence the number of detected lensed events. The first term describes a purely geometrical effect of how the comoving volume at a given redshift varies with a change in cosmological parameters. The second term describes the change in the distribution of lenses due to changes in the structure formation. Third shows how the fractional area covered by lenses at a given redshift varies due to the geometric effect.
@@ 2.2.2 Expected distribution of lensing time delays:
  • The redshift distribution of GW sources: We expect that this can be measured with sufficient precision from the large number $\sim 10^{6}$ of unlensed events that will dominate the data (see, e.g., [42]). In section 3 we forecast the prospective constraints on cosmological parameters assuming various theoretical models of the source redshift distribution (figure 1). In section 4 we study how uncertainties and errors in inferring the source redshift distribution can affect the constraints on cosmological parameters (figure 2).
@@ 3 Effect of source distribution models
Figure 1: Left: Different models for the distributions of redshifts for BBH mergers. These include prediction from population synthesis studies by Dominik [47] and Belczynski [16, 49], as well as uniform in comoving volume. Other models are based on Madau-Dickinson star formation rate [50] and consider different distributions of the time delay between formation and merger. Middle: The distributions of time delay between two lensed images of a source for different models of redshift distribution of BBH mergers shown in the left panel. Dashed lines represent the actual time delay distributions $p(\Delta t~{}|~{}\vec{\Omega})$ while solid lines represent time delay distributions $p(\Delta t~{}|~{}\vec{\Omega},T_{\mathrm{obs}})$ that will be observable in $10\;\mathrm{yrs}$. Time delay distributions are calculated for $\vec{\Omega}=\{\Omega_{m}=0.316,\;H_{0}=67.3\}$. Right: Posteriors (95% credible region) on cosmological parameters $\left(\Omega_{m},H_{0}\right)$ for different models of redshift distributions. This is done considering a BBH merger rate of $5\times 10^{5}\;\mathrm{yr}^{-1}$ with observation time period of $10\;\mathrm{yrs}$. Dashed cross represents true cosmological parameters.
The left panel of figure 1 displays the redshift distributions of BBH mergers which we consider. We consider a “true” cosmology $\vec{\Omega}_{\mathrm{true}}=\{\Omega_{m}=0.316,\;H_{0}=67.3\}$. The true distributions of lens redshift and parameters are calculated using a HMF model as described by [51], implemented in HMFcalc package [52] as described in section 2.2. We assume a merger rate of $R=5\times 10^{5}\;\mathrm{yr}^{-1}$ for BBHs, with an observation time period $T_{\mathrm{obs}}=10\;\mathrm{yrs}$.
@@
Using the method outlined in section 2.1, we compute the posteriors on the cosmological parameters $\vec{\Omega}$ from the different observing scenarios corresponding to the different source redshift distributions. We also assume that the redshift distribution of the sources is measured with sufficient precision from the observation of unlensed events. Posteriors for the cosmological parameters are shown in the right panel of figure 1. When the source populations extend to high redshift (e.g., Dominik, Uniform), the precision is better in comparison to populations where the merger redshift does not have support at high redshift (e.g., Belczynski, flatlog, etc.). This is because the lensing optical depth increases with redshift, and therefore, populations that extend to high redshift are expected to have a larger number of lensed events than the populations that don’t extend to high redshift. The expected 95% credible intervals in the posteriors of the cosmological parameters are summarised in table 1, for different source redshift distribution models.
4 Effect of errors in measuring the source redshift distribution
Figure 2: Left panel: The red curve shows the redshift distribution model based on the Madau-Dickinson star formation rate, with an exponential delay-time distribution with $\tau=0.1\;\mathrm{Gyr}$. Light cyan curves show the reconstructed (posterior predictive) distributions as shown in figure 5 of [42], while the solid cyan curve shows their average. Middle Panel: The 95% credible regions of the posteriors of $\Omega_{m}$ and $H_{0}$. The lensed events are simulated using the red curve in the left panel. The cosmological inference is done with the ‘true’ redshift distribution (red contour), as well as with 100 reconstructed source redshift distributions (light cyan contours, corresponding to the light cyan curves in the left panel). The cyan posteriors are biased due to the biased reconstruction of the source redshift distribution. Right Panel: This panel shows the effect of the marginalisation over the source redshift distribution uncertainties. Here, the lensed events are simulated using the solid cyan curve in the left panel, so that there is no systematic bias. The solid cyan contour shows the posterior estimated using the same redshift distribution. The dashed blue contour shows the posterior that is marginalised over the source redshift distribution uncertainties. The broadening of the posterior due to this effect is minimal.
In section 3, we assumed that the true redshift distribution of the sources is accurately known from the observation of unlensed events. However, this is not an entirely valid assumption as uncertainties and errors in the measurement of luminosity distance could bias our estimation of the redshift distribution, leading to systematic errors as well as additional statistical uncertainties in the estimation of cosmological parameters. Here we investigate the severeness of this effect.
Combining the inferred luminosity distance posteriors from a number of BBH events and assuming a cosmology, we will be able to reconstruct the redshift distribution of BBH mergers in the universe. This can be done either using a parametric model of the population properties of BBHs (such as their mass and redshift distribution) or using non-parametric methods [53]. The errors in the reconstruction of the redshift distribution of BBH mergers using XG detectors were studied in [42], using a parametric model. We use the results of their study for characterising the corresponding errors in cosmological inference. As done in [42], we use a ‘true’ redshift distribution given by a model based on Madau-Dickinson star formation rate and exponential delay time distribution with a time scale, $\tau=0.1\;\mathrm{Gyr}$ (red curves in figures 1 and 2).
We use the posterior predictive distributions of the merger rate density from  [42], which was derived from the posteriors on the parameters of the redshift distribution model (figures 4 and 5 of [42]). We take $\sim 100$ samples of the merger rate density distributions and convert them into source redshift distributions. In the left panel of figure 2, the light cyan curves represent these posterior predictive distributions while the solid cyan curve shows their average. These reflect the uncertainties in the measurement of the redshift distribution, while the red curve shows the ‘true’ distribution.
Now we simulate lensed events using the ‘true’ redshift distribution and infer cosmological parameters using different posterior predictive redshift distributions (corresponding to different light cyan lines in the left panel of figure 2). The middle panel of figure 2 shows the posteriors of the cosmological parameters inferred using the ‘true’ redshift distribution (red line) and 100 sampled redshift distributions (thin cyan lines). It is evident that there is a systematic bias in the recovery, which is attributed to the biased reconstruction of the merger rate density as seen in the left panel of figure 2.
We also explore a scenario in which the redshift distribution of BBHs is estimated without any systematic biases, but with some statistical uncertainty. In order to simulate such a scenario, we use the average of 100 sampled redshift distributions as the ‘true’ distribution (cyan curve in the left panel of figure 2). We then use this ‘true’ distribution to simulate lensed events and to infer cosmological parameters (right panel of figure 2). Here, the solid cyan contour represents the posterior (95% credible region) of cosmological parameters using the new ‘true’ redshift distribution, showing no systematic bias. When we factor in the uncertainties in the estimation of source redshift distribution making use of the posterior predictive distributions (light cyan curves in the left panel of  2), the posteriors of the cosmological parameters will have a scatter similar to the light cyan curves in the middle panel of figure 2. The dashed dark blue contour shows the posterior that is marginalised over the uncertainties in the estimation of the source redshift distribution. We can see that the broadening of the posterior is minimal.
In summary, the expected statistical uncertainties in the estimation of the source redshift distribution have a negligible effect on cosmological inference. However, the redshift distribution needs to be estimated without any systematic bias. In this preliminary investigation, we have neglected the correlation of the parameters of the source distribution model with the cosmological parameters. We anticipate that this oversight will not significantly broaden the posteriors on cosmological parameters, as the expected measurement error on the merger rate density is small [42]. Our future plans include conducting a comprehensive analysis that considers all the parameters of the source redshift distribution along with the cosmological parameters and marginalising over them.